### v1

@@ 4 Summary
The S-wave fully heavy ($QQ\bar{Q}\bar{Q}$), doubly heavy ($QQ^{(\prime)}\bar{q}\bar{q}$), and singly heavy ($Qs\bar{q}\bar{q}$) tetraquark systems are investigated within the quark potential model. The complex scaling method is used to identify bound and resonant states, and the Gaussian expansion method is used to solve the complex-scaled Schrödinger equation. Good candidates for experimental states in different systems are found, including fully charmed tetraquark resonances $X(6900),X(7200)$ [5, 6, 7], doubly charmed tetraquark bound state $T_{cc}(3875)^{+}$ [1, 2], and charm-strange tetraquark resonance $T_{\bar{c}\bar{s}0}^{*}(2870)$ [3, 4]. Moreover, a series of tetraquark states are predicted and may be searched for in near future experiments, such as $bc\bar{q}\bar{q},cs\bar{q}\bar{q},bs\bar{q}\bar{q}$ bound states, $bc\bar{b}\bar{c},cc\bar{q}\bar{q},bc\bar{q}\bar{q},bs\bar{q}\bar{q}$ resonances, and so on. The spatial structures of tetraquark states are identified by the rms radii. We find that all fully heavy tetraquarks have compact tetraquark configuration, all low-lying heavy-strange tetraquark are meson molecules, while doubly heavy tetraquarks exhibit rich configurations, including meson molecules, compact even tetraquarks (QCD hydrogen molecule), compact diquark-antidiquark tetraquarks, and compact diquark-centered tetraquarks (QCD helium atom). The various internal structures reveal different binding mechanisms of tetraquark states.

### v2

@@ 4 Summary
The S-wave fully heavy ($QQ\bar{Q}\bar{Q}$), doubly heavy ($QQ^{(\prime)}\bar{q}\bar{q}$), and singly heavy ($Qs\bar{q}\bar{q}$) tetraquark systems are investigated within the quark potential model. The complex scaling method is used to identify bound and resonant states, and the Gaussian expansion method is used to solve the complex-scaled Schrödinger equation. Good candidates for experimental states in different systems are found, including fully charmed tetraquark resonances $X(6900),X(7200)$ LHCb:2020bwg , ATLAS:2023bft , CMS:2023owd , doubly charmed tetraquark bound state $T_{cc}(3875)^{+}$ LHCb:2021vvq , LHCb:2021auc , and charm-strange tetraquark resonance $T_{\bar{c}\bar{s}0}^{*}(2870)$ LHCb:2020bls , LHCb:2020pxc . Moreover, a series of tetraquark states are predicted and may be searched for in near future experiments, such as $bc\bar{q}\bar{q},cs\bar{q}\bar{q},bs\bar{q}\bar{q}$ bound states, $bc\bar{b}\bar{c},cc\bar{q}\bar{q},bc\bar{q}\bar{q},bs\bar{q}\bar{q}$ resonances, and so on. The spatial structures of tetraquark states are identified by the rms radii. We find that all fully heavy tetraquarks have compact tetraquark configuration, all low-lying heavy-strange tetraquark are meson molecules, while doubly heavy tetraquarks exhibit rich configurations, including meson molecules, compact even tetraquarks (QCD hydrogen molecule), compact diquark-antidiquark tetraquarks, and compact diquark-centered tetraquarks (QCD helium atom). The various internal structures reveal different binding mechanisms of tetraquark states.